A sail is in the form of a right triangle that is three times as high as it is wide. The sail is made from 6 square meters of material. What is the height?
a=1/2bh

Answer by Cromlix(4381)

Width = x
Height = 3x
Area = 1/2 (base * height)
6m^2 = 1/2 (x * 3x)
6m^2 = 1/2 3x^2
Cross multiply.
12m^2 = 3x^2
4m^2 = x^2
x = 2m
Height = 3x
Height = 6m
Hope this helps.
:-)